Find two consective even integers such that the smaller added to three times the larger gives a sum of 54. the numbers are.?

Grade:12

2 Answers

Vikas TU
9 years ago
Your smaller number, let's call n. The other number is a consecutive even number, so you add 2. So, n and 2 + n are your two integers. Equation: 
n + 3( 2 + n ) = 54 
Distribute the 3: 
n + 6 + 3n = 54 
Add the like terms ( n and 2n ): 
4n + 6 = 54 
Subtract 6 from both sides: 
4n = 48 
Divide by 4 on both sides: 
n = 12 
So if n = 12, then n + 2 = 14. 
*Your numbers are 12 and 14*
Rickey
9 years ago
thnx!